OpenStudy (anonymous):

Which of the following is equivalent to log 8 35 rounded to three decimal places

OpenStudy (anonymous):

1.710 1.588 0.641 0.585

OpenStudy (anonymous):

\[\log_8(35)=\frac{\ln(35)}{\ln(8)}\] and a calculator

OpenStudy (anonymous):

is there a way to do this by hand

OpenStudy (precal):

no, you would have to use tables of logs. They stop teaching that method in the mid 80s. satellite is using change of base (best way to do it now).

OpenStudy (precal):

you could do it by hand, if the solution was a whole number example but for your problem, you will need a calculator